Titration of 0.2121 g of pure Na2C2O4 (134.00 g/mol) required 43.31 mL

of KMnO4.What is the molar concentration of the KMnO4 solution?


1
Expert's answer
2022-05-09T17:03:10-0400

Solution:

Balanced chemical equation:

5Na2C2O4 + 2KMnO4 + 8H2SO4 → 2MnSO4 + 5Na2SO4 + K2SO4 + 8H2O + 10CO2

According to stoichiometry:

Moles of Na2C2O4 / 5 = Moles of KMnO4 / 2

or:

2 × (Mass of Na2C2O4 / Molar mass of Na2C2O4) = 5 × (Molarity of KMnO4 × Volume of KMnO4)

Molarity of KMnO4 = (2 × Mass of Na2C2O4) / (5 × Volume of KMnO4 × Molar mass of Na2C2O4)

Molarity of KMnO4 = (2 × 0.2121 g) / (5 × 0.04331 L × 134.00 g mol−1) = 0.01462 mol/L

Molarity of KMnO4 = 0.01462 M


Answer: The molar concentration of the KMnO4 solution is 0.01462 M